I need the slope for 4x-4=0, I can not find my notes for the section...
Answer by Alan3354(69443)

4x-4=0
x = 1
------
For any value of y, x = 1 is a line parallel to the y-axis.
The slope is infinite, or undefined.
----------------
0y = x - 1
y = x/0 - 1/0
m = 1/0
b = -1/0
--> undefined